# Why Fine-Tuning Struggles with Forgetting in Machine Unlearning? Theoretical Insights and a Remedial Approach

### v1

@@ 3.2 Overlapping Features
The datasets $\mathbf{X}_{f}$ and $\mathbf{X}_{r}$ possess $d_{r}$ overlapped features, while $\mathbf{w}_{*}$ embodies the coefficients applicable across all features.
Remark 2.
Under 2, the dataset can be structured as follows: $\mathbf{X}_{r}^{\top}=[\mathbf{R}^{\top},\mathbf{L}_{1}^{\top},\mathbf{0}]$ and $\mathbf{X}_{f}^{\top}=[\mathbf{0},\mathbf{L}_{2}^{\top},\mathbf{F}^{\top}]$, where $\mathbf{R}\subseteq\mathbb{R}^{d_{r}\times n_{r}}$ and $\mathbf{F}\subseteq\mathbb{R}^{d_{f}\times n_{f}}$ represent the distinct features of the remaining and forgetting data, respectively. $\mathbf{L}_{1}\subseteq\mathbb{R}^{d_{lap}\times n_{r}}$ and $\mathbf{L}_{2}\subseteq\mathbb{R}^{d_{lap}\times n_{f}}$ denote the overlapped parts. Similarly to the 1, $d_{r}$ and $d_{f}$ are the distinct feature numbers for remaining and forgetting data, respectively, while the equation $d_{r}+d_{lap}+d_{f}=d$ holds. Additionally, the optimal solution can be decomposed into $\mathbf{w}_{*}=\mathbf{w}_{*}^{f}+\mathbf{w}_{*}^{lap}+\mathbf{w}_{*}^{r}$ such that $\mathbf{y}^{f}=\mathbf{X}_{f}^{\top}(\mathbf{w}_{*}^{f}+\mathbf{w}_{*}^{lap})$ and $\mathbf{y}^{r}=\mathbf{X}_{r}^{\top}(\mathbf{w}_{*}^{r}+\mathbf{w}_{*}^{lap})$.
@@
Theorem 2 shows that the previous conclusions remain valid under the assumptions of overlapping features, as the information from all training data, including forget data, is preserved from the pretrained model, $\mathbf{w}_{o}$, to the unlearned model through fine-tuning, $\mathbf{w}_{t}$. Consequently, the loss on both the remaining dataset and the forgetting dataset for the fine-tuning model is zero. Additionally, an interesting observation is that the number of overlapping features does not impact the unlearning accuracy of the fine-tuning model.
Figure 1: Machine Unlearning Performance via (Regularized) Fine-tuning with (without) Overlapping Features. Figure 1(a) and Figure 1(b) present the relationship between machine unlearning loss (i.e. RA, UA) and the number of fine-tuning data samples under distinct features and overlapping features assumptions, using naive FT method. In contrast, Figure 1(c) and Figure 1(d) show the same relationship using regularized fine-tuning methods, as discussed in Section 4.
@@ 4 Eliminating Forgetting Data Features from Pre-Trained Model Enhances Unlearning
  • 1.
    Distinct Features Scenario. When the features of $D_{r}$ and $D_{f}$ are distinct, we construct $\hat{\mathbf{w}}_{o}$ by retaining only the components corresponding to $D_{r}$ and setting the rest to zero. Formally, we define $\hat{\mathbf{w}}_{o}$ as $\hat{\mathbf{w}}_{o}^{\prime}[0:d_{r}]=\mathbf{w}_{o}[0:d_{r}]$ or equivalently can be understood as:
@@
  • 2.
    Overlapping Features Scenario. When features overlap across $D_{r}$ and $D_{f}$, we consider two cases:
@@
      Option A (Retaining Overlapping Features): We retain the overlapping features between $D_{r}$ and $D_{f}$, which can be expressed as $\hat{\mathbf{w}}_{o}[0:d_{r}+d_{o}]=\mathbf{w}_{o}[0:d_{r}+d_{o}]$ or equivalently
      $\hat{\mathbf{w}}_{o}[i]=\begin{cases}\mathbf{w}_{o}[i],&\text{if }i\in\text{features of }D_{r}\cup\text{overlapping features},\\ 0,&\text{otherwise}.\end{cases}$
@@
Let $\mathbf{w}_{o}$ be a pretrained model obtained the overall dataset $D=D_{r}\cup D_{f}$. Before performing unlearning (fine-tuning), we modify $\mathbf{w}_{o}$ to remove the components associated with $D_{f}$ as described above. Then, using the modified models $\hat{\mathbf{w}}_{o}$ ($\hat{\mathbf{w}}_{o}^{\prime}$) in the unlearning process, we have:
@@
    Distinct Features Scenario (Under the 1), we have:
    RL: $L(\hat{\mathbf{w}_{t}},D_{r})=0$; UL: $L(\hat{\mathbf{w}_{t}},D_{f})=\|\mathbf{w}_{*}^{f}\|^{2}_{\frac{1}{n_{f}}\mathbf{X}_{f}\mathbf{X}_{f}^{\top}}$,
@@
    Overlapping Features Scenario (Under 2), we have:
@@
According to Theorem 3, under the distinct features assumption, the regularized unlearned model achieves the same remaining and unlearning loss as the golden model. Furthermore, when considering overlapping features, if the overlapped component from the pretrained model is retained, the remaining loss remains zero, as with the golden model, while the unlearning loss differs only in the projection component. This difference can be considered negligible when applied to $\mathbf{w}_{*}^{r}$ and $\mathbf{w}_{*}^{lap}$ due to the model assumption. Figure 1(c) and Figure 1(d) verify our theoretical conclusions. However, if the overlapped component is discarded from the pretrained model, the remaining loss is no longer zero, and there is a small change to the unlearning loss that can be overlooked. These findings offer several insights into the design of machine unlearning algorithms: 1) Regularization on the pretrained model can significantly improve unlearning accuracy while preserving the remaining accuracy. If we can identify the component of the pretrained model related to the forgetting data, applying regularization to this component can further enhance UA. Our theorem also explains recent related works, such as [17, 6], which apply a mask to the pretrained model either randomly or by regularizing the weights associated with the forgetting data to provide better unlearn performance. These methods share the same underlying principle discussed here.
@@
Figure 2: Comparison of Machine Unlearning Loss with and without Overlapping Features. Figure 2(a) retains overlapping features from the pretrained model, showing the matching performance between regularized $\mathbf{w}_{t}$ model and golden model $\mathbf{w}_{g}$; Figure 2(b) discards the overlapping features, showing a decline in remaining accuracy.
2) When considering overlapping features, retaining them does not substantially affect unlearning accuracy, but discarding them compromises the remaining accuracy. As shown in Theorem 3, the remaining loss can not retain zero unless the remaining data $\mathbf{X}_{r}$ can be fully represented by the fine-tuning space, meaning $\mathbf{P}_{t}\mathbf{X}_{r}=\mathbf{X}_{r}$. Additionally, as the number of overlapping features increases, the impact on both remaining and unlearning loss becomes more significant. Discarding too many overlapping components can lead to instability in the remaining accuracy, as the model loses essential information needed to represent $D_{r}$, which in turn causes the remaining loss to increase. Figure 1 and Figure 2 validate our theoretical findings.
5 From Theory to Method: Discriminative Regularization
Figure 3: Performance comparison of fine-tuning methods on CIFAR-10 and CIFAR-100 datasets using five metrics: Unlearning Accuracy (UA), MIA-Efficacy, Retaining Accuracy (RA), Test Accuracy (TA), and Run Time. Each metric is normalized to the range [0, 1] based on the best result across all unlearning methods for ease of visualization, with the actual best value provided alongside each metric.
In Section 4 we show that once the components of the pretrained model related to the forgetting data are identified and removed, unlearning accuracy can be significantly improved. However, in practice, the training dataset and model are often not well-structured (1 and 2 may not hold). In such cases, as motivated by Equation 5, we know that the fine-tuning space may fail to unlearn from the forgetting data, allowing all information from the pretrained model to be retained. This raises the question: what happens if the fine-tuning space learns incorrect or faulty information about the forgetting data? To ensure that fine-tuning not only preserves the utility of the model on remaining data but also effectively forgets the target data, we introduce a regularized constraint on the fine-tuning unlearn process. Specifically, we encourage the model to forget by learning incorrect labels for the target data, which leads the following optimization problem:
[…]
where the first term is the cross entropy loss function used to measure the model’s accuracy on the remaining dataset; $\alpha$ is a regularization parameter that balances the trade-off between maintaining accuracy on the remaining data and forgetting the target data; the second term corresponds to the Kullback–Leibler divergence function applied to the target data, which is intentionally mislabeled [8]. This term acts as a penalty, encouraging the model to reduce its ability to accurately predict the target data, thereby facilitating the unlearning process.
Comparisons with prior work. We compare our approach with related work on regularized fine-tuning methods. One recent work [6] similar to ours is based on Saliency-based unlearning by minimizing the following optimization problem:
[…]
The primary difference between Equation 6 and Equation 7 lies in two keys: $\mathbf{1)}$ Type of Loss Functions. Equation 6 employs a combination of cross-entropy loss to maintain accuracy and KL-divergence to control forget accuracy, emphasizing aligning the mislabeled model’s predictions with the target data distribution. In contrast, Equation 7 relies exclusively on cross-entropy loss for both retain and forget accuracy, thus focusing more directly on minimizing classification errors. $\mathbf{2)}$ Regularization Focus. Equation 6 primarily aims to preserve retaining performance with regularization to penalize distributional differences. On the other hand, Equation 7 focuses on enhancing forget accuracy with regularization to retain the remaining accuracy. Notably, the regularization parameter is typically constrained to the range $(0,1]$.
However, based on our previous analysis, we favor the principle that regularization should prioritize remaining accuracy over unlearning accuracy (since retaining overlapping features does not significantly impact UA, but discarding them compromises RA). Therefore, we will explore the impact of switching the regularization focus in Equation 7, which we refer to as Inverse CE (ICE).
6 Experiment

### v2

@@ 3.2 Overlapping Features
The datasets $\mathbf{X}_{f}$ and $\mathbf{X}_{r}$ possess $d_{r}$ overlapped features, which can be structured as follows: $\mathbf{X}_{r}^{\top}=[\mathbf{R}^{\top},\mathbf{L}_{1}^{\top},\mathbf{0}]$ and $\mathbf{X}_{f}^{\top}=[\mathbf{0},\mathbf{L}_{2}^{\top},\mathbf{F}^{\top}]$, where $\mathbf{R}\subseteq\mathbb{R}^{d_{r}\times n_{r}}$ and $\mathbf{F}\subseteq\mathbb{R}^{d_{f}\times n_{f}}$ represent the distinct features of the remaining and forgetting data, respectively. $\mathbf{L}_{1}\subseteq\mathbb{R}^{d_{lap}\times n_{r}}$ and $\mathbf{L}_{2}\subseteq\mathbb{R}^{d_{lap}\times n_{f}}$ denote the overlapped parts.
Similarly to 1, $d_{r}$ and $d_{f}$ are the distinct feature numbers for remaining and forgetting data, while the equation $d_{r}+d_{lap}+d_{f}=d$ holds. It can also indicates that the optimal solution can be decomposed into $\mathbf{w}_{*}=\mathbf{w}_{*}^{f}+\mathbf{w}_{*}^{lap}+\mathbf{w}_{*}^{r}$ such that $\mathbf{y}^{f}=\mathbf{X}_{f}^{\top}(\mathbf{w}_{*}^{f}+\mathbf{w}_{*}^{lap})$ and $\mathbf{y}^{r}=\mathbf{X}_{r}^{\top}(\mathbf{w}_{*}^{r}+\mathbf{w}_{*}^{lap})$.
@@
Theorem 2 shows that the previous conclusions remain valid under the assumptions of overlapping features, as the information from all training data, including forget data, is preserved from the pretrained model, $\mathbf{w}_{o}$, to the unlearned model through fine-tuning, $\mathbf{w}_{t}$. Consequently, the loss on both the remaining dataset and the forgetting dataset for the fine-tuning model is zero. Additionally, an interesting observation is that the number of overlapping features does not impact the unlearning accuracy of the fine-tuning model. The proof of Theorem 2 is provided in Section B.3.
@@ 4 Eliminating Forgetting Data Features from Pre-Trained Model Enhances Unlearning
1. Distinct Features Scenario. When the features of $D_{r}$ and $D_{f}$ are distinct, we construct $\hat{\mathbf{w}}_{o}$ by retaining only the components corresponding to $D_{r}$ and setting the rest to zero. Formally, we define $\hat{\mathbf{w}}_{o}$ as $\hat{\mathbf{w}}_{o}^{\prime}[0:d_{r}]=\mathbf{w}_{o}[0:d_{r}]$ or equivalently can be understood as:
@@
2. Overlapping Features Scenario. When features overlap across $D_{r}$ and $D_{f}$, we consider two cases:
@@
    Option A (Retaining Overlapping Features): We retain the overlapping features between $D_{r}$ and $D_{f}$, which can be expressed as $\hat{\mathbf{w}}_{o}[0:d_{r}+d_{lap}]=\mathbf{w}_{o}[0:d_{r}+d_{lap}]$ or equivalently
    $\hat{\mathbf{w}}_{o}[i]=\begin{cases}\mathbf{w}_{o}[i],&\text{if }i\in\text{features of }\\ &D_{r}\cup\text{overlapping features},\\ 0,&\text{otherwise}.\end{cases}$
@@
Let $\mathbf{w}_{o}$ be a pretrained model obtained the overall dataset $D=D_{r}\cup D_{f}$. Before performing unlearning (fine-tuning), we modify $\mathbf{w}_{o}$ to remove the components associated with $D_{f}$ as described above. Then, using the modified models $\hat{\mathbf{w}}_{o}$ ($\hat{\mathbf{w}}_{o}^{\prime}$) in the unlearning process yields
@@
    In the distinct features scenario (i.e., under the 1), we have:
    RL: $L(\hat{\mathbf{w}_{t}},D_{r})=0$; UL: $L(\hat{\mathbf{w}_{t}},D_{f})=\|\mathbf{w}_{*}^{f}\|^{2}_{\frac{\mathbf{X}_{f}\mathbf{X}_{f}^{\top}}{n_{f}}}$,
@@
    In the overlapping features scenario (i.e., under 2), we have:
@@
According to Theorem 3, under the distinct features assumption, the masked unlearned model achieves the same remaining and unlearning loss as the golden model. Furthermore, when considering overlapping features, if the overlapped component from the pretrained model is retained, the remaining loss remains zero, as with the golden model, while the unlearning loss differs only in the projection component. This difference can be considered negligible when applied to $\mathbf{w}_{*}^{r}$ and $\mathbf{w}_{*}^{lap}$ due to the model assumption. Figure 1(c) and Figure 1(d) verify our theoretical conclusions. However, if the overlapped component is discarded from the pretrained model, the remaining loss is no longer zero, and there is a small change to the unlearning loss that can be overlooked. These findings offer several insights into the design of machine unlearning algorithms: 1) Masking on the pretrained model can significantly improve unlearning accuracy while preserving the retaining accuracy. If we can identify the component of the pretrained model related to the forgetting data, applying masking to this component can further enhance UA. Our theorem also explains recent related works, such as [17, 6], which apply a mask to the pretrained model either randomly or by regularizing the weights associated with the forgetting data to provide better unlearn performance. These methods share the same underlying principle discussed here. However, one overlook a critical scenario: when the remaining data and forgetting data share similar features, it becomes unclear whether those shared features should be retained. As shown in Theorem 3, our work provides a clear resolution to this question. 2) When considering overlapping features, retaining them does not substantially affect unlearning accuracy, but discarding them compromises the retaining accuracy. As shown in Theorem 3, the remaining loss can not retain zero unless the remaining data $\mathbf{X}_{r}$ can be fully represented by the fine-tuning space, meaning $\mathbf{P}_{t}\mathbf{X}_{r}=\mathbf{X}_{r}$. Additionally, as the number of overlapping features increases, the impact on both remaining and unlearning loss becomes more significant. Discarding too many overlapping components can lead to instability in the retaining accuracy, as the model loses essential information needed to represent $D_{r}$, which in turn causes the remaining loss to increase. Figure 1 and Figure 2 validate our theoretical findings. The proof of Theorem 3 is provided in Section B.4.
@@
Figure 2: Comparison of Machine Unlearning Loss with and without Overlapping Features. Figure 2(a) retains overlapping features from the pretrained model, showing the matching performance between masked $\mathbf{w}_{t}$ model and golden model $\mathbf{w}_{g}$; Figure 2(b) discards the overlapping features, showing a decline in retaining accuracy.
5 Rethinking Masking in Approximate Unlearning
5.1 Discarding Overlapping Features May Harm Retaining Accuracy
In Section 4 we show that once the components of the pretrained model related to the forgetting data are identified and removed, unlearning accuracy can be significantly improved. In practice, an existing work [6] shares the same principles as discussed above. Specifically, it constructs the desired weight saliency map by leveraging the gradient of a loss function with respect to the model weights on the forgetting dataset:
$\mathbf{m}_{f}=\mathbbm{1}\left(\left|\nabla_{\mathbf{w}}L\left(\mathbf{w};{D}_{f}\right)\right|_{\mathbf{w}=\mathbf{w}_{{o}}}\mid\geq\gamma\right),$ (6)
where $\mathbbm{1}(g\geq\gamma)$ is an element-wise indicator function that outputs 1 for the $i$-th element if $g_{i}\geq\gamma$ and 00 otherwise, ||$\cdot$|| ⋅ | denotes the element-wise absolute value operation, and $\gamma>0$ is a hard threshold. Variables with larger gradients, which are highly associated with the forgetting dataset, are identified and subsequently masked. Then, the unlearn model can be updated by:
$\mathbf{w}_{u_{f}}=\underbrace{\mathbf{m}_{f}\odot(\Delta\mathbf{w}+\mathbf{w}_{o})}_{\text{salient weights }}+\underbrace{(\mathbf{1}-\mathbf{m}_{f})\odot\mathbf{w}_{o}}_{\text{original weights }},$ (7)
where $\mathbf{w}_{o}$ represents the original model, $\mathbf{1}$ denotes an all-one vector and $\Delta\mathbf{w}$ is optimized by the following problem:
$\underset{\Delta\mathbf{w}}{\min}\mathcal{L}(\mathbf{w}_{u_{f}}):=$ $\mathbb{E}_{(\mathbf{x},y)\sim{D}_{f},y^{\prime}\neq y}[\ell_{\mathrm{CE}}(\mathbf{w}_{u_{f}};\mathbf{x},y^{\prime})]+\alpha\mathbb{E}_{(\mathbf{x},y)\sim{D}_{r}}[\ell_{\mathrm{CE}}(\mathbf{w}_{u_{f}};\mathbf{x},y)],$ (8)
where $y^{\prime}$ is the random label of the image $\mathbf{x}$ different from $y$, $\alpha$ is a regularization parameter and $\ell_{\mathrm{CE}}$ is the cross-entropy (CE) loss for supervised classification.
However, Equation 7 overlooks the scenario where the remaining data and the forgetting data share similar model weights, making it unclear whether those weights should be preserved. In Theorem 3, we demonstrate that when considering overlapping features (features shared between the remaining dataset and the forgetting dataset), retaining them is more beneficial than discarding them, which contrasts slightly with the approach in [6]. Specifically, we show that discarding overlapping features compromises retaining accuracy while retaining them ensures better performance for the remaining dataset.
Based on this analysis, we propose an alternative principle–Retention-Based Masking: the weight saliency map should be constructed based on the remaining dataset rather than the forgetting dataset:
$\mathbf{1}-\mathbf{m}_{r}=\mathbf{1}-\mathbbm{1}\left(\left|\nabla_{\mathbf{w}}L\left(\mathbf{w};{D}_{r}\right)\right|_{\mathbf{w}=\mathbf{w}_{{o}}}\mid\geq\gamma\right).$ (9)
Therefore, the unlearn model will be updated by
$\mathbf{w}_{u_{r}}=\underbrace{(\mathbf{1}-\mathbf{m}_{r})\odot(\Delta\mathbf{w}+\mathbf{w}_{o})}_{\text{salient weights }}+\underbrace{\mathbf{m}_{r}\odot\mathbf{w}_{o}}_{\text{original weights }}.$ (10)
Here, $\Delta\mathbf{w}$ is optimized using Equation 8 with the retention-based mask, which combines the forgetting loss on the random labels of forgotten data with the retaining loss on the remaining data.
Table 2: Cifar-10, Cifar-100 Comparison. We evaluate the unlearning performance on Cifar datasets using Dense models only. Evaluation metrics from Section 5.2 are employed. Metrics are reported as mean $\pm$ standard deviation across five seeds. (UA: Unlearning Acc., RA: Retaining Acc., TA: Test Acc.)
Cifar-10 Random Forgetting
Methods UA MIA-Efficacy RA TA Avg. Disparity Run Time (min)
Retrain $5.80_{\pm 0.12}$ $13.91_{\pm 0.15}$ $100.00_{\pm 0.00}$ $94.30_{\pm 0.13}$ $0.00$ $82.15$
FT $0.18_{\pm 0.04}(5.62)$ $1.70_{\pm 0.10}(12.21)$ $99.92_{\pm 0.07}(0.08)$ $94.25_{\pm 0.15}(0.05)$ $4.48$ $2.51$
SalUn $2.38_{\pm 0.31}(3.42)$ $15.45_{\pm 0.14}(1.54)$ $99.62_{\pm 0.22}(0.38)$ $94.11_{\pm 0.23}(0.19)$ $1.39$ $\mathbf{2.50}$
Ours $3.24_{\pm 0.12}(\mathbf{2.56})$ $14.22_{\pm 0.11}(\mathbf{0.31})$ $99.72_{\pm 0.20}({0.28})$ $93.90_{\pm 0.10}({0.40})$ $\mathbf{0.88}$ $\mathbf{2.50}$
Cifar-10 Class-wise Forgetting
Methods UA MIA-Efficacy RA TA Avg. Disparity Run Time (min)
Retrain $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $94.81_{\pm 0.09}$ $0.00$ $82.00$
FT $8.36_{\pm 3.03}(91.64)$ $40.76_{\pm 8.03}(59.24)$ $99.92_{\pm 0.03}(\mathbf{0.08})$ $94.41_{\pm 0.29}(0.40)$ $37.84$ $2.50$
SalUn $99.78_{\pm 0.15}(0.22)$ $100.00_{\pm 0.00}(\mathbf{0.00})$ $99.47_{\pm 0.23}(0.53)$ $93.55_{\pm 0.44}({1.26})$ $0.50$ $\mathbf{2.50}$
Ours $99.98_{\pm 0.02}(\mathbf{0.02})$ $100.00_{\pm 0.00}(\mathbf{0.00})$ $99.69_{\pm 0.30}({0.31})$ $93.44_{\pm 0.30}(1.37)$ $\mathbf{0.42}$ $\mathbf{2.50}$
Cifar-100 Random Forgetting
Methods UA MIA-Efficacy RA TA Avg. Disparity Run Time (min.)
Retrain $24.75_{\pm 0.11}$ $49.68_{\pm 0.35}$ $99.98_{\pm 0.01}$ $74.57_{\pm 0.06}$ 0.00 82.33
FT $0.11_{\pm 0.03}(24.64)$ $5.66_{\pm 0.47}(44.02)$ $99.97_{\pm 0.01}(\mathbf{0.01})$ $75.45_{\pm 0.17}(\mathbf{0.88})$ 16.65 2.50
SalUn $27.75_{\pm 0.33}(3.00)$ $71.42_{\pm 0.11}(21.74)$ $98.65_{\pm 0.29}(1.33)$ $68.97_{\pm 0.39}(5.60)$ $7.91$ $\mathbf{1.3}$
Ours $23.53_{\pm 0.22}(\mathbf{1.22})$ $69.02_{\pm 0.09}(\mathbf{19.34})$ $98.74_{\pm 0.12}(1.24)$ $69.16_{\pm 0.11}(5.41)$ $\mathbf{6.80}$ $\mathbf{1.3}$
Cifar-100 Class-wise Forgetting
Methods UA MIA-Efficacy RA TA Avg. Disparity Run Time (min)
Retrain $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $99.98_{\pm 0.01}$ $73.75_{\pm 0.20}$ $0.00$ $82.22$
FT $11.55_{\pm 7.1}(88.45)$ $59.33_{\pm 17.57}(40.67)$ $99.78_{\pm 0.03}(0.20)$ $74.61_{\pm 0.30}(0.86)$ $32.55$ $2.52$
SalUn $100.00_{\pm 0.00}(\mathbf{0.00})$ $100.00_{\pm 0.00}(\mathbf{0.00})$ $99.63_{\pm 0.10}(0.35)$ $74.76_{\pm 0.31}(1.01)$ $0.35$ $\mathbf{2.50}$
Ours $100.00_{\pm 0.00}(\mathbf{0.00})$ $100.00_{\pm 0.00}(\mathbf{0.00})$ $99.90_{\pm 0.06}\mathbf{({0.08})}$ $74.47_{\pm 0.72}(\mathbf{0.72})$ $\mathbf{0.19}$ $\mathbf{2.50}$
5.2 Experiment